# Scaling Laws Are Unreliable for Downstream Tasks: A Reality Check

### v1

@@ Abstract
Downstream scaling laws aim to predict task performance at larger scales from pretraining losses at smaller scales. Whether this prediction should be possible is unclear: some works demonstrate that task performance follows clear linear scaling trends under transformation, whereas others point out fundamental challenges to downstream scaling laws, such as emergence and inverse scaling. In this work, we conduct a meta-analysis of existing data on downstream scaling laws, finding that close fit to linear scaling laws only occurs in a minority of cases: 39% of the time. Furthermore, seemingly benign changes to the experimental setting can completely change the scaling trend. Our analysis underscores the need to understand the conditions under which scaling laws succeed. To fully model the relationship between pretraining loss and downstream task performance, we must embrace the cases in which scaling behavior deviates from linear trends.
@@
Nicholas Lourie1thanks: Equal contribution.  Michael Y. Hu [^1] footnotemark: 1  Kyunghyun Cho1,2,3 1New York University  2Prescient Design  3 CIFAR LMB {nick.lourie, michael.hu, kyunghyun.cho}@nyu.edu
@@ 1 Introduction
Scaling laws for language models establish that pretraining validation loss or perplexity improves reliably when increasing the size of the model, compute, and training data Kaplan et al. (2020); Hoffmann et al. (2022); Pearce and Song (2024). However, better pretraining loss does not always translate to better downstream performance (Magnusson et al., 2024). This gap can be caused by a variety of issues: among the best known is emergence, or the fact that on some tasks, models below a certain parameter scale show no trend or near-chance performance Wei et al. (2022). Furthermore, model performance can increase and then decrease, in what is called inverse scaling McKenzie et al. (2023); Wilcox et al. (2024). Despite these challenges, several works also claim that downstream performance is roughly linear with respect to simple transformations of the pretraining loss Huang et al. (2024); Gadre et al. (2024); Chen et al. (2025).
@@
Thus, the extent to which downstream scaling laws work is unclear. How can downstream scaling laws follow linear forms when we know of many tasks that exhibit emergence or inverse scaling? Are these phenomena edge cases, or are they prevalent, with no established explanation? Here, we aim to clarify this confusion. We identify and consider three core factors affecting downstream scaling laws: 1) the data used for pretraining and validation, 2) the downstream task, and 3) the experimental setup. Realistic changes to each of these factors change the relationship to downstream performance, to the point that the scaling law’s functional form might no longer hold.
We find that:
@@
    Choosing a different dataset for computing validation perplexity can flip scaling trends. For example, models trained on one pretraining corpus might appear to improve faster than another pretraining corpus, but this trend can reverse when using a different validation dataset (§3).
@@
    Examining a previous study (Gadre et al., 2024), we find that predictable scaling only occurs in a minority of cases: 39% of the time. This indicates that phenomena like emergence and inverse scaling are actually quite common, and can represent a significant proportion of existing downstream tasks (§4).
@@
    The scaling behavior observed under one experimental setting may not hold for other settings: a task with predictable scaling in one setting might show nonmonotonic scaling or even no trend at all in another setting (§5).
Our analysis suggests that scaling laws are context-specific. As such, we cannot assume downstream scaling will always be strictly linear. Rather, we need to better understand the failure modes of existing scaling laws and develop a holistic (and perhaps more complex) model of how models improve on downstream tasks.
@@ 2 Background
Unlike pretraining, there is little consensus on how to approach downstream scaling laws. Early efforts attempted to predict downstream performance directly from parameter count, data, or compute (Ivgi et al., 2022; Mahmood et al., 2022; OpenAI, 2023), but downstream performance often showed a noisy relationship to these quantities (Tay et al., 2022). Other efforts sought surrogates for scale, such as inferred latent variables (Ruan et al., 2024) or task-specific losses (Grattafiori et al., 2024; Bhagia et al., 2024), in hopes of mapping from compute to the surrogate, and from the surrogate to downstream performance.
Many works have found that downstream scaling laws are more stable when stated in terms of pretraining loss (Xia et al., 2023; Huang et al., 2024; Gadre et al., 2024; Du et al., 2024; Chen et al., 2025), the most widely used surrogate for scale. In particular, if two models have a different number of parameters or see a different number of tokens but attain the same pretraining loss, then they tend to achieve the same downstream performance (Xia et al., 2023; Gadre et al., 2024; Du et al., 2024). So pretraining loss correlates with downstream performance to a surprising degree; however, as we will show, this relationship is far from absolute.
In the best case, downstream task performance is roughly linear in some monotonic transformation of validation loss, e.g.
$y=a\exp\{c\cdot x\}+b,$
where $y$ is the task performance metric and $x$ is the validation perplexity Gadre et al. (2024).
Linearity is amenable to extrapolation, the ultimate goal of downstream scaling laws, but linearity is known not to hold in all cases. In cases with emergent capabilities or U-shaped scaling, results at small scales cannot be used to predict capabilities at larger scales, because there exists a structural break Caballero et al. (2023): the function that describes one part of the scaling curve does not apply globally, and thus one cannot extrapolate from small to large. In these cases, one cannot make the global assumption of linearity under a monontonic transformation. Emergence can sometimes be mitigated by choosing a different downstream metric Schaeffer et al. (2023), but in other cases breakthrough improvements remain “stubbornly emergent” Zhao et al. (2025). In general, downstream performance can behave nonmonotonically as a function of scale Gadre et al. (2024); Magnusson et al. (2025). Li et al. (2025) found scaling laws generally difficult to reproduce and sensitive to the scaling law’s functional form, training setup, data collection, and fitting algorithm.
@@ 3 Scaling Laws Are Specific to the Data
In this section, we show that downstream scaling laws depend on three factors: 1) pretraining data, 2) validation data, and 3) the downstream task. It is unclear how we should choose the validation data, and yet its choice can entirely reverse which pretraining setup appears superior for a downstream task. Choosing a validation dataset is necessary to compare models pretrained on different data, as the pretraining losses are otherwise incomparable because they are taken over different distributions.
Here, we examine the evaluation results from Gadre et al. (2024), who pretrained models over different corpora. We observe how computing validation perplexity on either C4 or Paloma’s 100 Programming Languages Raffel et al. (2020); Magnusson et al. (2024) changes downstream scaling trends on the CoQA and HellaSwag tasks Reddy et al. (2019); Zellers et al. (2019).
@@
For the HellaSwag task with C4 as the validation corpus (top left), pretraining on either C4 or RedPajama produces the same scaling law. However, when using 100 Programming Languages (100 PLs, top right), the scaling laws for C4 and RedPajama no longer superimpose—pretraining on C4 appears to achieve much better performance even for a worse validation loss.
@@ 4 Irregular Scaling Is Common
Phenomena like emergence and inverse scaling suggest that linear scaling laws do not capture all task scaling behavior. However, it is also unclear how prevalent these phenomena are in practice. Thus, we re-examine scaling behavior on the 46 tasks tested by Gadre et al. (2024), classifying them into six categories visualized in Figure 2. We find that linear scaling actually occurs in a minority of cases in their setting: 39% of the time (Figure 1). For some experimental setups, non-linear scaling is actually the norm. Are the experimental choices of Gadre et al. (2024) abnormal? On the contrary, all three of pretraining corpora, validation datasets, and downstream tasks are from well-known sources in the literature. Downstream tasks are comprised of established evaluations like BoolQ, HellaSwag, and BIG-Bench Clark et al. (2019); Zellers et al. (2019); Srivastava et al. (2023). Irregular scaling occurs within popular tasks and is easy to find.
@@ 5 Scaling Behavior Is Not Always Robust
Finally, we show that conclusions about scaling laws may not generalize across settings: setups with the same validation data and downstream tasks may observe entirely different scaling trends. To show this, we take the 10 overlapping tasks between Gadre et al. (2024) and Magnusson et al. (2025). These authors consider several of the same pretraining corpora and downstream tasks; however, their implementation details differ (see Appendix D). For example, Gadre et al. (2024) use fewer answer choices for Commonsense QA. Thus, we might expect to see quantitative differences in some of their scaling laws, but—what is more surprising—is that we also see qualitative changes in their scaling behavior.
To plot downstream performance on the same axes, we evaluate all models from Gadre et al. (2024) and Magnusson et al. (2025) on the same validation corpus (C4). For Magnusson et al. (2025), which does not use C4, we evaluate 200 released checkpoints, which vary in their pretraining setup: [^1] The 20M models are the smallest models in Magnusson et al. (2025) and have the highest cross-entropy loss.
@@
    Pretraining corpus: {Dolma, C4, DCLM-Baseline, RefinedWeb (Falcon), FineWeb-Pro} Soldaini et al. (2024); Raffel et al. (2020); Li et al. (2024); Penedo et al. (2023b); Zhou et al. (2025).
@@
Out of the 10 overlapping downstream tasks, Figure 4 shows three that produce different scaling trends between setups. In both setups, MMLU shows a positive trend; however, the noisiness and shape of that trend differs greatly. At an even greater extreme, CommonsenseQA’s scaling behavior qualitatively changes: while CommonsenseQA shows nonmonotonic scaling in Gadre et al.’s (2024) results, it exhibits a clean scaling law with Magnusson et al.’s (2025). On the other hand, in Magnusson et al.’s (2025) results, BoolQ’s scaling law appears trendless; however, this lack of trend comes from the models spanning too small a validation loss range. Since Gadre et al.’s (2024) models cover a wider range of validation losses, the trend more clearly emerges. Thus, scaling behavior can fluctuate, even between controlled studies on pretraining corpora.
@@ 6 Discussion
As practitioners of scaling laws, we must realize that predictable scaling laws often exist, but one cannot assume that they hold for all contexts. Even if scaling is stable on the same task for the same validation data, other aspects of the experimental setup might change the scaling behavior (§5). To some extent, scaling laws are investigator-specific Li et al. (2025), and so each investigator must verify the scaling law’s presence with visualizations and regression diagnostics Shalizi (2015).
For researchers, downstream scaling laws offer many fascinating new directions. Empirically, we need better ways to stabilize scaling laws and detect when irregular scaling might occur. We must understand the factors affecting scaling laws, and what parts of the experimental setup must remain constant for linear scaling laws to hold. Theoretically, we need a model for why predictable scaling occurs Hutter (2021), and a core goal of such a theory should be explaining exactly the cases in which scaling is unpredictable.
@@ 7 Conclusion
Our work uses data and model checkpoints from existing studies Gadre et al. (2024); Magnusson et al. (2025). While this is sufficient for the counterexamples featured in this work, there may be unknown biases shared between these two projects that we have missed.
@@ Acknowledgments
Appendix A Scaling Behavior Examples
Figures 5 through 10 each presents all the tasks studied in Gadre et al. (2024) that demonstrated a specific scaling behavior.
@@ Appendix A Scaling Behavior Examples
Figure 10: Tasks in Gadre et al. (2024) where scaling behavior is better described by a sigmoid functional form.
Appendix B Sensitivity to Experimental Setups
@@
Appendix C The Effect of the Pretraining and Validation Corpora
Figure 12 compares how the task, the pretraining, and validation corpora affect the scaling curve. For ease of visualization, we show only the effect on the tasks with the cleanest, most predictable scaling laws.
@@ Appendix C The Effect of the Pretraining and Validation Corpora
Appendix D Data Sources
Finally, LLM Foundry’s version of SIQA had an error where the gold labels were incorrect. [^6] https://github.com/mosaicml/llm-foundry/pull/774. This issue was fixed in LLM Foundry v0.5.0; [^7] https://github.com/mosaicml/llm-foundry/releases/tag/v0.5.0 however, Gadre et al. (2024) used v0.4.0 for at least some of their experiments. [^8] https://wandb.ai/samir/dcnlp/runs/rezso5ec/files/requirements.txt As a result, we do not examine SIQA in our analyses, although Figure 11 in Appendix B includes SIQA for completeness.
Appendix E Reproducibility

### v2

@@ Abstract
Downstream scaling laws aim to predict task performance at larger scales from the model’s performance at smaller scales. Whether such prediction should be possible is unclear: some works discover clear linear scaling trends after simple transformations of the performance metric, whereas others point out fundamental challenges to downstream scaling laws, such as emergence and inverse scaling. In this work, we conduct a meta-analysis of existing data on downstream scaling laws, and we find that predictable scaling only occurs in a minority of cases: 39% of the time. Moreover, seemingly benign changes to the experimental setting can completely change the scaling behavior. Our analysis underscores the need to understand the conditions under which scaling laws succeed. To accurately model the relationship between pretraining loss and task performance, we must embrace the cases in which scaling behavior deviates from linear trends.
@@
Nicholas Lourie1thanks: Equal contribution. Code: https://github.com/nicholaslourie/scale-fails   Michael Y. Hu [^1] footnotemark: 1   Kyunghyun Cho1,2,3 1New York University  2Prescient Design  3 CIFAR LMB {nick.lourie, michael.hu, kyunghyun.cho}@nyu.edu
@@ 1 Introduction
Scaling laws for pretraining establish that the loss improves reliably when increasing the size of the model, training data, or compute (Kaplan et al., 2020; Hoffmann et al., 2022; Pearce and Song, 2024). However, better pretraining loss does not always translate to better downstream performance (Magnusson et al., 2024). This gap can be caused by a variety of issues; among the best known is emergence, or the fact that on some tasks models below a certain scale show no trend or near-chance performance (Wei et al., 2022). In addition, model performance can increase then decrease, in what is called inverse scaling (McKenzie et al., 2023; Wilcox et al., 2024). Despite these challenges, several works also find downstream performance is roughly linear in the pretraining loss, possibly after a transformation (Huang et al., 2024; Gadre et al., 2025; Chen et al., 2025).
@@
Thus, the extent to which downstream scaling laws work is unclear. How can they follow linear forms when we know of many tasks that exhibit emergence or inverse scaling? Are these just edge cases, or are they common, with no established explanation? Here, we aim to clarify this confusion. We explore and consider three core factors affecting downstream scaling laws: 1) the data used for pretraining and validation, 2) the downstream task, and 3) the experimental setup. Realistic changes to each of these factors can change the relationship to downstream performance, to the point that the scaling law’s functional form might no longer hold. The behavior can even change qualitatively—with a decrease in trend under one set of conditions flipping to an increase under another. Such qualitative changes can not be removed by a new functional form. They pose fundamental obstacles to studying large scale models with small scale proxies.
In particular, we find that:
@@
    Choosing a different dataset for the validation perplexity can flip scaling trends. For instance, models pretrained on one corpus might appear to improve faster than another, but this trend can reverse with different validation data (§3).
@@
    Revisiting a prior study (Gadre et al., 2025), we only find predictable scaling in a minority of cases: 39%. Phenomena like emergence or inverse scaling can actually be quite common, occurring for many downstream tasks (§4).
@@
    Scaling behavior from one experimental setup can qualitatively change under another; a task with predictable scaling could become non-monotonic or even show no trend at all (§5).
Our analysis suggests that scaling laws are context-specific. As such, we cannot assume downstream scaling will always be strictly linear. Rather, we need to better understand the failure modes of existing scaling laws and develop a holistic (and perhaps more complex) model of how foundation models improve on downstream tasks.
@@ 2 Background
Unlike pretraining, there is little consensus on how to approach downstream scaling laws. Early efforts tried to predict downstream performance directly from parameters, data, or compute (Ivgi et al., 2022; Mahmood et al., 2022; OpenAI, 2023), but downstream performance often showed a noisy relationship to these quantities (Tay et al., 2022). Other efforts sought surrogates for scale, like latent capabilities (Ruan et al., 2024) or task-specific losses (Grattafiori et al., 2024; Bhagia et al., 2024), in the hope of mapping from compute to surrogate, and from surrogate to downstream performance.
Many works have found downstream scaling laws are more stable when stated in terms of pretraining loss, the most widely used surrogate for scale (Xia et al., 2023; Huang et al., 2024; Du et al., 2024; Gadre et al., 2025; Chen et al., 2025). Specifically, if two models differ in their number of parameters or pretraining tokens but still attain the same pretraining loss, then they tend to achieve the same downstream performance (Xia et al., 2023; Du et al., 2024; Gadre et al., 2025). Some authors generalize this principle, viewing pretraining loss as a general way to compare models—not just with different scales but different training recipes (Huang et al., 2024; Chen et al., 2025). However, while pretraining loss correlates with downstream performance to a surprising degree, this relationship is far from absolute (Tay et al., 2022).
In the best case, downstream task performance is roughly linear in some monotonic transformation of validation cross-entropy loss ($x$) and downstream metric ($y$):
f(y)=ag(x)+bf(y)=a\,g(x)+b
For instance, Gadre et al. (2025) relate error rate $y$ to the validation loss $x$ via: $y=\epsilon-k\,\exp\{-\gamma\,x\}$. The functions, $f$ and $g$, must depend on the metric; in some cases, a linear relationship alone could be enough (Huang et al., 2024; Chen et al., 2025).
Such global structure enables extrapolation—the ultimate goal of downstream scaling laws; however, predictable structure might not exist. Phenomena such as emergence, inverse, and U-shaped scaling preclude this kind of global structure (Wei et al., 2022; McKenzie et al., 2023; Wei et al., 2023). They destroy it by creating structural breaks, or points where the function describing one part of the scaling curve does not describe another (Andrews, 1993). Without global structure, it is impossible to extrapolate from small to large. Certain tasks, like multiple choice questions, are more susceptible to emergence (Schaeffer et al., 2025). Sometimes emergence can be mitigated by choosing a different downstream metric (Schaeffer et al., 2023), but in other cases breakthrough improvements remain “stubbornly emergent” (Zhao et al., 2025; Du et al., 2024). Li et al. (2025) found scaling laws generally difficult to reproduce and sensitive to the functional form, training setup, data collection, and fitting algorithm, while Magnusson et al. (2025) found that comparing models at a smaller scale—without extrapolation—performed as well or better than scaling laws in determining the best data mix.
@@ 3 Scaling Laws Are Specific to the Data
Downstream scaling laws depend on several factors, including the pretraining data, validation data, and downstream task. If you vary the pretraining data, then you must fix a validation corpus to compare the loss across models. Once it is fixed, you might hope to focus on the validation loss alone and simplify your research. It is unclear how to choose the correct validation data but, what is more, its choice can entirely reverse which pretraining setup appears superior for a downstream task if we do not consider the full context. The pretraining data, validation data, and downstream task all interact in forming the scaling law—you can not consider one without the other two. In particular, you can not compare data mixes by their validation losses or downstream scaling laws based upon them.
To evince these claims, we reexamine the results from Gadre et al. (2025), who pretrained models over different corpora. We observe how pretraining data, validation data, and task all interact in forming the scaling law (Figure 3). The colors correspond to pretraining corpora: C4 (Raffel et al., 2020) and RedPajama (Weber et al., 2024); the columns correspond to validation data: C4 (Raffel et al., 2020) and Paloma’s 100 Programming Languages (Magnusson et al., 2024); and the rows correspond to tasks: CoQA (Reddy et al., 2019) and HellaSwag (Zellers et al., 2019). If you ignore the other factors, then choice of validation data can completely change the scaling trend:
@@
For HellaSwag with C4 as the validation corpus (top left), pretraining on either C4 or RedPajama produces the same scaling law. However, when using 100 Programming Languages (100 PLs, top right), the scaling laws for C4 and RedPajama no longer superimpose—pretraining on C4 appears to achieve much better performance even for a worse validation loss.
@@ 4 Irregular Scaling Is Common
Phenomena like emergence and inverse scaling suggest that linear scaling laws do not capture all task scaling behavior. However, it is also unclear how prevalent these phenomena are in practice. Thus, we re-examine scaling behavior on the 46 tasks tested by Gadre et al. (2025), classifying them into six categories visualized in Figure 2. We find that linear scaling actually occurs in a minority of cases in their setting: 39% of the time (Figure 1). For some experimental setups, non-linear scaling is actually the norm. Are the experimental choices of Gadre et al. (2025) abnormal? On the contrary, all three of pretraining corpora, validation datasets, and downstream tasks are from well-known sources in the literature. Downstream tasks are comprised of established evaluations like BoolQ, HellaSwag, and BIG-Bench (Clark et al., 2019; Zellers et al., 2019; Srivastava et al., 2023). Irregular scaling occurs within popular tasks and is easy to find.
@@ 5 Scaling Behavior Is Not Always Robust
Finally, we show that conclusions about scaling laws may not generalize across settings: setups with the same validation data and downstream tasks may observe entirely different scaling trends. To show this, we take the 10 overlapping tasks between Gadre et al. (2025) and Magnusson et al. (2025). These authors consider several of the same pretraining corpora and downstream tasks; however, their implementation details differ (see Appendix A). For example, Gadre et al. (2025) use fewer answer choices for Commonsense QA. Thus, we might expect to see quantitative differences in some of their scaling laws, but what is more surprising is that we also see qualitative changes in their scaling behavior.
To enable comparison, we evaluate all models from Gadre et al. (2025) and Magnusson et al. (2025) on the same validation corpus (C4). For Magnusson et al. (2025), which does not use C4, we evaluate 200 released checkpoints, which vary in their pretraining setup: [^1] The 20M models are the smallest models in Magnusson et al. (2025) and have the highest cross-entropy loss.
@@
    Pretraining corpus: {Dolma, C4, DCLM-Baseline, RefinedWeb (Falcon), FineWeb-Pro} (Soldaini et al., 2024; Raffel et al., 2020; Li et al., 2024; Penedo et al., 2023b; Zhou et al., 2025).
@@
Out of the 10 overlapping downstream tasks, Figure 4 shows three that produce different scaling trends between setups. In both setups, MMLU shows a positive trend; however, the noisiness and shape of that trend differs greatly. At an even greater extreme, CommonsenseQA’s scaling behavior qualitatively changes: while CommonsenseQA shows nonmonotonic scaling in Gadre et al.’s (2025) results, it exhibits a clean scaling law with Magnusson et al.’s (2025). On the other hand, in Magnusson et al.’s (2025) results, BoolQ’s scaling law appears trendless; however, this lack of trend comes from the models spanning too small a validation loss range. Since Gadre et al.’s (2025) models cover a wider range of validation losses, the trend more clearly emerges. Thus, scaling behavior can fluctuate, even between controlled studies for the same task with the same pretraining corpora.
@@ 6 Discussion
As practitioners of scaling laws, we must realize that predictable scaling laws often exist, but one cannot assume that they hold for all contexts. Even if scaling is stable on the same task for the same validation data, other aspects of the experimental setup might change the scaling behavior (§5). To some extent, scaling laws are investigator-specific (Li et al., 2025), and so each investigator must verify the scaling law’s presence with visualizations and regression diagnostics (Shalizi, 2015).
For researchers, downstream scaling laws offer many fascinating new directions. Empirically, we need better ways to stabilize scaling laws and detect when irregular scaling might occur. We must understand the factors affecting scaling laws, and what parts of the experimental setup must remain constant for linear scaling laws to hold. Theoretically, we need a model for why predictable scaling occurs (Hutter, 2021), and a core goal of such a theory should be explaining exactly the cases in which it does not.
@@ 7 Conclusion
Our work uses data and model checkpoints from existing studies (Gadre et al., 2025; Magnusson et al., 2025). While this is sufficient for the counterexamples featured in this work, there may be unknown biases shared between these two projects that we have missed.
@@ Acknowledgments
Appendix A Data Sources
Finally, LLM Foundry’s version of SIQA had an error where the gold labels were incorrect. [^6] https://github.com/mosaicml/llm-foundry/pull/774. This issue was fixed in LLM Foundry v0.5.0; [^7] https://github.com/mosaicml/llm-foundry/releases/tag/v0.5.0 however, Gadre et al. (2025) used v0.4.0 for at least some of their experiments. [^8] https://wandb.ai/samir/dcnlp/runs/rezso5ec/files/requirements.txt As a result, we do not examine SIQA in our analyses, although Figure 11 in Appendix D includes SIQA for completeness.
Appendix B Reproducibility
Appendix C Scaling Behaviors
Our analysis centers on qualitative scaling behavior in order to avoid having its conclusions depend on a particular form of the scaling law. Without formal criteria, researchers might disagree on particular examples, but then researchers might also disagree on formal criteria. Thus we define the scaling behaviors informally, as follows:
predictable
Performance increases, without too much variation around the trend.
inverse
Performance decreases as loss improves.
nonmonotonic
Performance switches between increasing and decreasing.
noisy
Performance increases but varies greatly around the trend.
trendless
Performance is flat or there is too much noise to discern a trend.
breakthrough
Performance starts flat, abruptly increases, and then plateaus.
Figures 5 through 10 present the tasks studied in Gadre et al. (2025) sorted by scaling behavior. The figures include scaling laws fitted using the functional form from Gadre et al. (2025) that relates the error rate to the validation loss: $y=\epsilon-k\,\exp\{-\gamma\,x\}$. We apply this functional form to the accuracy using the fact that accuracy is one minus the error rate. The scaling laws were fitted by minimizing mean squared error with differential_evoluation from SciPy (Virtanen et al., 2020). For the optimization, we used bounds of 0 to 1 for $\epsilon$, 0 to 20 for $\ln(k)$ (we searched $k$ on a log scale to widen the range), and 0 to 20 for $\gamma$. In addition, we used 60 for the popsize multiplier. Figures 5 through 10 also include the $R^{2}$ of each scaling law’s fit as a measure of how closely the scaling behavior adheres to the functional form.
Appendix D Sensitivity to Experimental Setups
Appendix E The Effect of the Pretraining and Validation Corpora
Figure 12 compares how the task, pretraining, and validation corpora affect the scaling curve. For ease of visualization, we only show the effect on tasks with the cleanest, most predictable scaling laws.
@@ Appendix E The Effect of the Pretraining and Validation Corpora
Figure 10: Tasks in Gadre et al. (2025) demonstrating breakthrough scaling.